# Mutatis Mutandis: Revisiting the Comparator in Discrimination Testing

### v2

@@ Abstract.
Testing for discrimination consists of deriving a profile, known as the comparator, similar to the profile making the discrimination claim, known as the complainant, and comparing the outcomes of these two profiles. An important aspect for establishing discrimination is evidence, often obtained via discrimination testing tools that implement the complainant-comparator pair. In this work, we revisit the role of the comparator in discrimination testing. We argue for the causal modeling nature of deriving the comparator, and introduce a two-kinds classification for the comparator: the ceteris paribus (CP), and mutatis mutandis (MM) comparators. The CP comparator is the standard one among discrimination testing, representing an idealized comparison as it aims for having a complainant-comparator pair that only differs on membership to the protected attribute. As an alternative to it, we define the MM comparator, which requires that the comparator represents what would have been of the complainant without the effects of the protected attribute on the non-protected attributes. The complainant-comparator pair, in that case, may also be dissimilar in terms of all attributes. We illustrate these two comparators and their impact on discrimination testing using a real illustrative example. Importantly, we position generative models and, overall, machine learning methods as useful tools for constructing the MM comparator and, in turn, achieving more complex and realistic comparisons when testing for discrimination.
algorithmic fairness, discrimination discovery, counterfactual model of discrimination
@@ 1. Introduction
The concept of discrimination refers to an unjustified difference in treatment toward an individual or group of individuals based on (perceived) membership to a protected-by-law group (Romei and Ruggieri, 2014). We view, thus, discrimination as unfairness sanctioned by law. Testing for discrimination against an individual instance is a comparative process centered around the complainant-comparator pair. As the names would suggest, the complainant is the individual instance motivating the discrimination claim in court, while the comparator is the individual instance used for validating the claim. We use the comparator to answer the hypothetical question behind any discrimination claim: would the decision have been the same if the complainant had been a member of the non-protected group instead of the protected group under non-discrimination law? The comparison between the outcomes of the complainant and its comparator helps establish whether the decision-making process in question is or not discriminatory. Most tools for discrimination testing, from the traditional methods like correspondence studies (Bertrand and Mullainathan, 2004) and situation testing (Bendick, 2007) to the recent algorithmic methods like discrimination discovery (Ruggieri et al., 2010) and counterfactual fairness (Kusner et al., 2017), rely on this pair. Importantly, these tools not only implement the complainant-comparator comparison but also derive the complainant’s comparator as a whole.
@@
Non-discrimination law requires similar, or similarly situated, individuals to be treated similarly regardless of gender, race, or other protected attributes. It is a requirement largely shared across countries, only with some variation on the list of protected attributes (Romei and Ruggieri, 2014). The comparator is meant to represent a similar (or, at least, comparable) individual to the complainant, allowing to envision an alternative path for the complainant conditional on “changing” its protected attribute. It requires what we view, in practice, as an imaginative act in which we find (e.g., via matching models (Qureshi et al., 2020)) or create (e.g., via generative models (Black et al., 2020)) an individual instance representative enough of the complainant’s alternative world. In fact, the comparator, unlike the complainant, can also be a fictitious individual instance (Weerts et al., 2023). In a time of modern decision flows in which machine learning (ML) methods are increasingly used to aid as well as regulate decision-making processes, such modeling flexibility around the comparator offers a promising ML application.
@@
The choice of comparator, meaning how these tools for discrimination testing derive the comparator, matters for three incremental reasons. First, it embodies a normative statement on what similarity between individuals is and, in turn, how the protected attribute influences individual outcomes within the context of interest (e.g., (Weerts et al., 2023)). Second, it conditions the results of these tools by determining which individual outcomes are to be compared for evaluating the discrimination claim (e.g., (Álvarez and Ruggieri, 2023)). Third, it acts as a target for the sort of substantive equality between individuals that we wish to achieve as a society through non-discrimination law (e.g., (Wachter et al., 2020)). Most tools aim to define the comparator as an observably similar instance to the complainant all but for membership to the protected attribute, such as defining a male candidate with the same work experience as the female complainant in a hiring decision or a white applicant with the same credit score as the non-white complainant in a lending decision. We refer to this standard operationalization of the comparator as the ceteris paribus, or “with all else being equal,” (CP) comparator. Precisely because of these reasons, debates around the CP comparator predate modern decision flows, e.g., (Westen, 1982; Heckman, 1998; Kohler-Hausmann, 2019), and remain relevant today as fair ML researchers tackle the new problem of algorithmic discrimination, e.g., (Kleinberg et al., 2019; Ruggieri et al., 2023; Álvarez et al., 2024). We would add that this reliance on the CP comparator is problematic for fair ML. It overlooks the potential positive role ML methods, particularly generative models, could play as tools for deriving more complex comparators that address the ongoing debates.
@@
This work studies the problem of testing for discrimination by revisiting the role of the comparator. We do so by, first, embracing the causal modeling nature of discrimination testing and, second, proposing an alternative to the CP comparator. Regarding the first point, given that the comparator answers to the “what would have been if” question around the complainant, we assert that the comparator is a counterfactual representation of the factual complainant. We further assert that the derivation of the comparator itself constitutes a causal modeling problem and that such problem formulation is, for now, the best possible one given the comparative nature of discrimination testing. Regarding the second point, we propose the mutatis mutandis, or “with the appropriate adjustments being made,” (MM) comparator. The MM comparator requires that all non-protected attributes constituting the comparator account for the downstream effects of the protected attribute. The CP comparator, thus, represents an idealized comparison to the complainant, while the MM comparator represents an adjusted comparison to the complainant. With the MM comparator, we expand the problem of discrimination testing and, given the greater complexity of deriving a MM comparator relative to a CP comparator (read, a stronger “imagination”), we identify an impactful application for fair ML models. Example 1.1 below motivates and illustrates the difference between the two comparators.
@@
(Tenure at U) This example is based on Morgan et al. (2021)’s research on the unequal impact of parenthood in academia for female researchers. [^1] Social expectations and household dynamics are key factors in creating this difference. Mothers, as the primary caregivers, tend to carry the burden of childbearing within their households, which consumes potential time that could be spent on advancing ones academic career by, e.g., writing papers at home or networking at conferences. Fathers tend to share less of this burden. See Morgan et al. (2021) for details. Also see Roman and Rodríguez (2023). Let us assume that the university U only considers the number of publications for granting tenure. Clara, a female researcher, was denied tenure for having published only 12 papers and files a discrimination claim against U, becoming the complainant. Suppose that we use Mike, a male researcher that also published 12 papers and was denied tenure by U, as her comparator. Do we deny Clara’s discrimination claim? Given what we know about parenthood’s impact on academic performance, would our answer change by considering that Clara is a mother? Assuming some quantifiable penalty on the number of publications from motherhood, would we be willing to use Vincent, with 18 publications and recently tenured, as Clara’s comparator instead of Mike?
@@
In Example 1.1 Mike represents the CP comparator (Definition 2.1) while Vincent the MM comparator (Definition 3.1). Depending on which one we use, assuming a literal comparison, we would invalidate or validate, respectively, Clara’s discrimination claim. Despite its simplicity, Example 1.1 captures the intricacies behind discrimination testing once we contextualize the decision-making process based on information about the protected attribute. Given what we know from works like Morgan et al. (2021), we are implicitly asking which male researcher is more similar to Clara? It is not an easy question to answer, highlighting the complexity behind modeling the comparator. With Example 1.1 and, overall, with this work we are not suggesting that the CP comparator is wrong or that the MM comparator is better. Instead, we want to show that the MM comparator is an alternative formulation to the CP comparator and, overall, stress the importance of the choice of comparator. Policies like non-discrimination law and positive actions aimed at addressing/removing systematic inequalities require concrete targets (Wachter et al., 2020). Discrimination testing tools, essentially, enforce such targets through the comparator, with the CP and MM comparators representing two distinct targets.
@@
In the remainder of this section we discuss our modeling choice for causality (Section 1.1) and relevant related work (Section 1.2). Moving forward, in Section 2 we argue that deriving the comparator is a causal modeling problem and formulate the CP comparator, which is the standard in discrimination testing. In Section 3, we propose the MM comparator and justify its departure from the CP comparator. In Section 4, using a popular real dataset in fair ML, we showcase both comparators by comparing two related discrimination testing tools—situation testing (Thanh et al., 2011) and counterfactual situation testing (Álvarez and Ruggieri, 2023)—that implement, respectively, the CP comparator and MM comparator. In Section 5, we provide a representative survey of discrimination testing tools. In Section 6, we conclude.
Contributions.
Our main contributions are threefold. First, we demonstrate why the derivation of the comparator is best understood as a causal modeling problem. Second, with the MM comparator we propose a new approach to deriving the comparator and, in turn, introduce the distinction between CP and MM comparators. Our experiments illustrate how this distinction impacts discrimination testing. Third, we position ML methods as a promising tool for achieving the MM comparator and developing better tools for discrimination testing.
Audience and scope.
@@
Discrimination testing is an interdisciplinary topic. Our work is aimed primarily at ML researchers working on this topic and, hence, our framing of it as a modeling problem. We are interested in quantitative tools used for deriving prima facie evidence for discrimination claims, which constitutes one aspect of the legal process for uncovering discrimination. Our focus is mainly on the operationalization of the comparator, which we view as a ML problem that holds for any modern decision flow, regardless of the nature of the decision-maker. Throughout the paper, we use generative model to refer to a model capable of creating synthetic data (McElreath, 2018). Similarly, we use counterfactuals to refer to interventional probability distributions obtained through causal reasoning (Pearl, 2009), not to be confused with counterfactual explanations222 Counterfactual explanations refer to individual instances similar to the one that is explained but have a different outcome. To avoid confusion, Miller (2019) proposes to name them contrastive explanations, and reserve the term counterfactual as understood in the statistical causality literature. from explainable AI (Stepin et al., 2021; Guidotti, 2024).
@@ 1.1. Auxiliary Causal Knowledge
We formulate causality using structural causal models (SCM) (Pearl, 2009), which is the most used framework within ML. SCM are probabilistic graphical models that link (latent) exogenous variables to (observable) endogenous variables via structural equations. Each SCM describes a probability distribution and can generate hypothetical distributions via interventions. Formally, a SCM $\mathcal{M}$ is a tuple $\mathcal{M}=\langle\mathbf{U},\mathbf{X},\mathbf{F}\rangle$, where $\mathbf{U}=(U_{1},\ldots,U_{p})$ are $p$ independent exogenous latent variables, with $\mathbf{U}\sim P(\mathbf{U})=\prod_{j}P(U_{j})$, $\mathbf{X}=(X_{1},\ldots,X_{p})$ are $p$ endogenous observed variables, and $\mathbf{F}=(f_{1},\ldots,f_{p})$ are structural equations such that:
(1) $\text{for $j=1,\dots,p$}\;\;\;X_{j}:=f_{j}(X_{pa(j)},U_{j})$
@@
Each $f_{j}$ defines the observed variable $X_{j}$ on the basis of its causal parents $X_{pa(j)}$ and of the latent variable $U_{j}$ Associated to the SCM $\mathcal{M}$ is the causal graph $\mathcal{G}$, in which each node is a random variable and the directed edges between them a causal effect. We assume a directed acyclical graph, or DAG, $\mathcal{G}$. Through $\mathcal{G}$ we represent how the variables relate to each other in terms of cause-effect pairs. For instance, the causal graph $X_{1}\rightarrow X_{2}$ reads as “$X_{1}$ causes $X_{2}$”.
@@
The SCM $\mathcal{M}$ in (1) can act as a generative model. In particular, it can generate counterfactual distributions that answer to counterfactual questions as the ones asked in discrimination claims. Let $A$ be one of the variables in $\mathbf{X}$ denoting the protected attribute. We calculate the counterfactual variables $\mathbf{X}^{CF}_{A\leftarrow\alpha}$ resulting from setting $A:=\alpha$ through a three-step procedure consisting of abduction, action, and prediction (Pearl et al., 2016). Abduction: for each prior distribution $P(U_{j})$ that describes $U_{j}$, we compute its posterior distribution given the evidence, or $P(U_{j}\;|\;\mathbf{X})$. Action: we intervene on $A$ by changing its structural equation to $A:=\alpha$, which gives way to a new SCM $\mathcal{M}^{\prime}$. Prediction: we generate the counterfactual distribution $P(\mathbf{X}^{CF}_{A\leftarrow\alpha}\;|\;\mathbf{X})$ by propagating the abducted $P(U_{j}\;|\;\mathbf{X})$ through the revised structural equations in $\mathcal{M}^{\prime}$.
@@
Beyond Section 4, in which we rely on counterfactual generation to implement the MM comparator, what we discuss in Sections 2 and 3 is not exclusive to SCM. It is our generative model of choice, mainly because previous works on discrimination testing (e.g., (Kusner et al., 2017; Kilbertus et al., 2017; Álvarez and Ruggieri, 2023)) use it. Here, our goal is not prioritize one generative model over the others, but to focus on one generative model to show how the others could be useful for the problem of discrimination testing.
1.2. Relevant Related Work

### v3

@@ Abstract.
Testing for individual discrimination involves deriving a profile, the comparator, similar to the one making the discrimination claim, the complainant, based on a protected attribute, such as race or gender, and comparing their decision outcomes. The complainant-comparator pair is central to discrimination testing. Most discrimination testing tools rely on this pair to establish evidence for discrimination. In this work we revisit the role of the comparator in discrimination testing. We first argue for the inherent causal modeling nature of deriving the comparator. We then introduce a two-kinds classification for the comparator: the ceteris paribus, or “with all else equal,” (CP) comparator and the mutatis mutandis, or “with the appropriate adjustments being made,” (MM) comparator. The CP comparator is the standard comparator, representing an idealized comparison for establishing discrimination as it aims for a complainant-comparator pair that only differs on membership to the protected attribute. As an alternative to it, we define the MM comparator, which requires that the comparator represents the “what would have been of” the complainant without the effects of the protected attribute on the non-protected attributes. Under the MM comparator, the complainant-comparator pair can be dissimilar in terms of the non-protected attributes, departing from an idealized comparison. Notably, the MM comparator is a more complex kind of comparator and its implementation offers an impactful venue for machine learning methods. We illustrate these two comparators and their impact on discrimination testing using a real-world example.
algorithmic fairness, discrimination discovery, counterfactual reasoning
@@ 1. Introduction
Discrimination refers to an unjustified difference in treatment toward an individual based on (perceived) membership to a protected by non-discrimination law group (Romei and Ruggieri, 2014). Testing for discrimination is a comparative process centered around the complainant-comparator pair. As the names would suggest, the complainant is the individual instance motivating the discrimination claim while the comparator is the individual instance used for validating it. We use the comparator to answer the hypothetical question behind all discrimination claims: would the decision have been the same if the complainant had been a member of the non-protected rather than the protected group? The comparison between the outcomes of the complainant and its comparator helps establish whether the decision-making process in question is or not discriminatory. Most tools for discrimination testing, from the traditional methods like correspondence studies (Bertrand and Mullainathan, 2004) and situation testing (Bendick, 2007) to the recent algorithmic methods like discrimination discovery (Ruggieri et al., 2010) and counterfactual fairness (Kusner et al., 2017), rely on this pair. These tools not only implement the complainant-comparator comparison but also derive the complainant’s comparator as a whole based on the relevant process data.
@@
The comparator is central to establishing evidence for discrimination. Non-discrimination law requires similar, or similarly situated, individuals to be treated similarly regardless of gender, race, or other protected attributes. It is a requirement shared across countries with only some variation on the list of protected attributes (Romei and Ruggieri, 2014). The comparator is meant to represent a similar individual to the complainant and allows to envision an alternative path for the complainant conditional on “changing” its protected attribute. It requires what we view in practice as an imaginative act in which we find (using, for instance, matching models (Qureshi et al., 2020)) or create (using generative models (Black et al., 2020)) an individual instance representative enough of the complainant’s alternative world. In fact, the comparator, unlike the complainant that motivates it, can be a fictitious individual instance (Weerts et al., 2023).
@@
The choice of comparator, meaning how these discrimination testing tools derive the comparator, matters for three incremental reasons. First, it embodies a normative statement on what similarity between individuals is and, in turn, how the protected attribute influences individual outcomes within the context of interest (Weerts et al., 2023). Second, it conditions the results of these tools by determining which individual outcomes are to be compared for evaluating the discrimination claim (Álvarez and Ruggieri, 2025). Third, it acts as a target for the sort of substantive equality between individuals that we wish to achieve as a society through non-discrimination law (Wachter et al., 2020). Most tools aim to define the comparator as an observably similar instance to the complainant all but for membership to the protected attribute, such as defining a male candidate with the same work experience as the female complainant in a hiring decision or a white applicant with the same credit score as the non-white complainant in a lending decision. We refer to this standard operationalization of the comparator as the ceteris paribus, or “with all else being equal,” (CP) comparator. Precisely because of these three reasons, debates around the CP comparator predate modern decision flows that rely on machine learning (ML) (Westen, 1982; Heckman, 1998; Kohler-Hausmann, 2019). These debates remain relevant as ML researchers tackle the new problem of algorithmic discrimination (Kleinberg et al., 2019; Ruggieri et al., 2023; Álvarez et al., 2024).
@@
This work studies the problem of testing for discrimination by revisiting the role of the comparator. We do so by, first, embracing the causal modeling nature of discrimination testing and, second, proposing an alternative to the standard CP comparator. Regarding the first point, given that the comparator answers to the “what would have been if” question around the complainant, we assert that the comparator is a counterfactual representation of the factual complainant. We further assert that the derivation of the comparator itself constitutes a causal modeling problem and that such problem formulation is the best possible one given the comparative nature of discrimination testing. Regarding the second point, we propose the mutatis mutandis, or “with the appropriate adjustments being made,” (MM) comparator. The MM comparator requires that all non-protected attributes constituting the comparator account for the downstream effects of the protected attribute. Hence, the CP comparator represents an idealized comparison to the complainant while the MM comparator represents an adjusted comparison to the complainant with each based on how we conceive modeling-wise the influence of the protected attribute on all other attributes. Example 1.1 motivates and illustrates the difference between the two comparators. Altogether, we view the implementation of the MM comparator as an impactful application for ML models. Implementing the MM comparator relative to the CP comparator imposes additional complexity (read, a stronger imaginative act) to the problem of discrimination testing that can be better addressed by ML models over more traditional models. This additional point contrasts with popular works, mainly within the fair ML field (Kohler-Hausmann, 2019; Hu and Kohler-Hausmann, 2020), that are adamant toward treating discrimination testing as a ML problem.
@@
(Tenure at U)  [^1] This example is based on Morgan et al. (2021)’s research on the unequal impact of parenthood in academia for female researchers. Social expectations and household dynamics are key factors in creating this difference. Mothers, as the primary caregivers, tend to carry the burden of childbearing within their households, which consumes potential time that could be spent on advancing ones academic career by, for instance, writing papers at home or networking at conferences. Fathers tend to share less of this burden. See Morgan et al. (2021) for details. Also see Roman and Rodríguez (2023) for similar results in couples outside of academia. Let us assume that the university U only considers the number of publications for granting tenure. Clara, a female researcher, was denied tenure for having published only 12 papers and files a discrimination claim against U, becoming the complainant. Suppose that we use Mike, a male researcher that also published 12 papers and was denied tenure by U, as her comparator. Do we deny Clara’s discrimination claim? Given what we know about parenthood’s impact on academic performance, would our answer change by considering that Clara is a mother? Assuming some quantifiable penalty on the number of publications from motherhood, would we be willing to use Vincent, with 18 publications and recently tenured, as Clara’s comparator instead of Mike?
@@
In Example 1.1, Mike represents the CP comparator (Definition 2.1) while Vincent represents the MM comparator (Definition 3.1). Depending on which one we use, assuming a literal comparison, we would invalidate or validate, respectively, Clara’s discrimination claim. Despite its simplicity, Example 1.1 captures the intricacies behind discrimination testing once we contextualize the decision-making process based on information about the protected attribute. Given what we know from works like Morgan et al. (2021), in this example we are implicitly asking which male researcher is more similar to Clara given what we know about gender and its role in this context? We believe that it is not an easy question to answer, highlighting the complexity behind modeling the comparator. With Example 1.1 and, overall, with this work we are not suggesting that the CP comparator is wrong or that the MM comparator is better. We simply want to position the MM comparator as a viable alternative formulation to the standard CP comparator and, in doing so, stress the importance of the choice of comparator in discrimination testing. Policies like non-discrimination law and positive actions aimed at addressing systematic inequalities require concrete targets (Wachter et al., 2020). Discrimination testing tools enforce such targets through the comparator. The CP and MM comparators, thus, represent two distinct targets. It is a missed opportunity to rely solely on the CP comparator, especially, when we already have ML modeling tools able to implement the more complex MM comparator. In this work, we focus on one of such ML tools by using structural causal models (SCM) Pearl (2009) to derive the MM comparator.
Contributions.
Our main contributions are twofold. First, we demonstrate why the derivation of the comparator is best understood as a causal modeling problem centered on counterfactual reasoning. This is indeed a divisive view on establishing discrimination but one that should be embraced for developing better discrimination testing tools. Second, with the MM comparator we propose a new approach to the derivation of the comparator and introduce the distinction between CP and MM comparators. This novel distinction expands the methodological landscape for establishing discrimination. Additionally, our illustrative experiment showcases how this distinction in comparators impacts discrimination testing in practice. The need for implementing the more imaginative MM comparator clearly positions ML models as promising approaches within discrimination testing.
@@
In the remainder of this section we discuss our modeling choice for causality (Section 1.1) and relevant related work (Section 1.2). In Section 2 we argue that deriving the comparator is a causal modeling problem and formulate the CP comparator. In Section 3, we propose the MM comparator and justify its departure from the CP comparator. In Section 4, using a popular real dataset in fair ML, we showcase both comparators by comparing two related discrimination testing tools—situation testing (Thanh et al., 2011) and counterfactual situation testing (Álvarez and Ruggieri, 2023, 2025)—that implement, respectively, the CP comparator and MM comparator. In Section 5, we provide a representative survey of discrimination testing tools. In Section 6, we conclude this work.
Audience, scope, and word choice.
@@
Discrimination testing is an interdisciplinary topic. Our work is aimed primarily at ML researchers. We are interested in the quantitative tools used for deriving prima facie evidence for discrimination claims, which constitutes one of many aspects of the legal process for establishing discrimination. [^2] Technically, all discrimination tools can only test for prima facie discrimination (Romei and Ruggieri, 2014; Weerts et al., 2023; Álvarez et al., 2024). This distinction is sometimes downplayed in the corresponding papers. We come back to this distinction in Section 2.2. Throughout the paper we focus on individual discrimination. Namely, we are interested at testing whether an individual instance of a protected group has been discriminated. The notion of group discrimination tackles instead discrimination against a subset of individuals from a protected group. [^3] However, Binns (2020) argues that the apparent conflict between individual and group fairness is an artifact of fairness metrics rather than reflecting different normative principles. Here, we stick with the current fair ML convention. Notably, although often used interchangeably, we make a clear distinction between unfair and discriminatory decisions. Both concepts focus on judging the decision outcomes around a specific attribute; however, discrimination has a stronger meaning than unfairness as it requires such attribute to be under non-discrimination law. Unfairness, in that sense, is more flexible in meaning. All discriminatory decisions are unfair, but not all unfair decisions are discriminatory. Discrimination, in short, can be viewed as unfairness sanctioned by law. Our focus is mainly on the operationalization of the comparator, which we view as a ML problem that holds for any modern decision flow, regardless of the nature of the decision-maker. Throughout the paper, we use generative model to refer to a model capable of creating synthetic data (McElreath, 2018). These are models capable of learning enough about the data generating process of interest to generate instances that are both novel and conceivable under said process, which includes, by definition, approaches like representation learning (Zemel et al., 2013) and SCM Pearl (2009). Similarly, we use counterfactuals to refer to interventional probability distributions obtained through causal reasoning (Pearl, 2009), not to be confused with counterfactual explanations [^4] Counterfactual explanations refer to individual instances similar to the one that is explained but have a different outcome. To avoid confusion, Miller (2019) proposes to name them contrastive explanations, and reserve the term counterfactual as understood in the statistical causality literature. from explainable AI (Stepin et al., 2021; Guidotti, 2024).
@@ 1.1. Auxiliary Causal Knowledge
We formulate causality using structural causal models (SCM) (Pearl, 2009), a popular framework within causal ML. SCM are probabilistic graphical models that link (latent) exogenous variables to (observable) endogenous variables via structural equations. Each SCM describes a probability distribution and can generate hypothetical distributions via interventions. Formally, a SCM $\mathcal{M}$ is a tuple $\mathcal{M}=\langle\mathbf{U},\mathbf{X},\mathbf{F}\rangle$, where $\mathbf{U}=(U_{1},\ldots,U_{p})$ are $p$ independent exogenous latent variables, with $\mathbf{U}\sim P(\mathbf{U})=\prod_{j}P(U_{j})$, $\mathbf{X}=(X_{1},\ldots,X_{p})$ are $p$ endogenous observed variables, and $\mathbf{F}=(f_{1},\ldots,f_{p})$ are structural equations such that:
(1) $\text{for $j=1,\dots,p$}\;\;\;X_{j}:=f_{j}(X_{pa(j)},U_{j}).$
@@
Each $f_{j}$ defines the observed variable $X_{j}$ on the basis of its causal parents $X_{pa(j)}$ and of the latent variable $U_{j}$ . Associated to the SCM $\mathcal{M}$ is the causal graph $\mathcal{G}$, in which each node is a random variable and the directed edges between them a causal effect. We assume a directed acyclical graph, or DAG, $\mathcal{G}$. Through $\mathcal{G}$ we represent how the variables relate to each other in terms of cause-effect pairs. The causal graph $X_{1}\rightarrow X_{2}$, for instance, reads as $X_{1}$ causes $X_{2}$”.
@@
The SCM $\mathcal{M}$ in (1) acts as a generative model by generating counterfactual distributions that answer to counterfactual questions as the ones asked in discrimination claims. Let $A$ be one of the variables in $\mathbf{X}$ denoting the protected attribute. We calculate the counterfactual variables $\mathbf{X}^{CF}_{A\leftarrow\alpha}$ resulting from setting $A:=\alpha$ through a three-step procedure consisting of abduction, action, and prediction (Pearl et al., 2016). Abduction: for each prior distribution $P(U_{j})$ that describes $U_{j}$, we compute its posterior distribution given the evidence, or $P(U_{j}\;|\;\mathbf{X})$. Action: we intervene on $A$ by changing its structural equation to $A:=\alpha$, which gives way to a new SCM $\mathcal{M}^{\prime}$. Prediction: we generate the counterfactual distribution $P(\mathbf{X}^{CF}_{A\leftarrow\alpha}\;|\;\mathbf{X})$ by propagating the abducted $P(U_{j}\;|\;\mathbf{X})$ through the revised structural equations in $\mathcal{M}^{\prime}$.
@@
Importantly, beyond Section 4, in which we rely on counterfactual generation to implement the MM comparator, what we discuss in Sections 2 and 3 is not exclusive to SCM. It is our generative model of choice, mainly because previous works on discrimination testing (Kusner et al., 2017; Kilbertus et al., 2017; Álvarez and Ruggieri, 2023, 2025) use it. Our goal is not prioritize one generative modeling approach over the others, but to focus on one generative model approach to show how the others could be useful for the problem of discrimination testing under the MM comparator.
1.2. Related Work